Exercise 25. Show that the exponential distribution for time to failure has the orgetfulness property: (6.5)
Answer #1

P( T \ge t0 + t1 | T \ge t0 )

= \frac{P( T \ge t_0 + t_1 \cap T \ge t_0 ) }{P(T \ge t_0)}

Now, the cdf of exponential distriution is,

P[T \ge t] = e^{-\lambda t} , t > 0

So, \frac{P( T \ge t_0 + t_1 \cap T \ge t_0 ) }{P(T \ge t_0)}

= \frac{P( T \ge t_0 + t_1 ) }{P(T \ge t_0)}

= \frac{ e^{-\lambda (t_1+t_0)} }{e^{-\lambda t_0}}

= e^{-\lambda t_1}

= P(T \ge t1) [Proved]